# Likelihood-Ratio Regularized Quantile Regression: Adapting Conformal Prediction to High-Dimensional Covariate Shifts

### v1

@@ Appendix H Lipschitz process
By Lemma L.3 and Lemma L.4, a global minimizer of the objective in Equation LR-QR exists, and since $\beta_{\textnormal{min}}\leqslant\beta_{\textnormal{lower}}$, $\beta_{\textnormal{max}}\geqslant\beta_{\textnormal{upper}}$, and $B\geqslant B_{\textnormal{upper}}$, any such minimizer lies in the interior of $\mathcal{H}_{B}\times\mathcal{I}$. Thus we may apply Lemma K.1 to the objective function $L$. We utilize two directional derivatives in the space $\mathcal{H}\times\mathbb{R}$. The first is in the direction $0_{\mathcal{H}}\times 1$, the unit vector in the $\beta$ coordinate. Since $(\hat{h},\hat{\beta})\in\mathcal{H}_{B}\times\mathcal{I}$, the magnitude of the second derivative of $L$ along this direction is bounded by $\nu_{1}$.
The second is in the direction of the vector $r_{\mathcal{H}_{B}}\times 0$, where $r_{\mathcal{H}_{B}}$ the projection of $r$ onto the closed convex set $\mathcal{H}_{B}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$. Since $(\hat{h},\hat{\beta})\in\mathcal{H}_{B}\times\mathcal{I}$, the magnitude of the second derivative of $L$ along this direction is bounded by $\nu_{2}$.
Given $\hat{h}$, let $\widehat{\textnormal{Cover}}(X):=\mathbb{P}\left[S\leqslant\hat{h}(X)|X\right]-(1-\alpha)$. Now, on the event $E$ that $L(\hat{h},\hat{\beta})-L(h^{*},\beta^{*})\leqslant\mathcal{E}_{\textnormal{gen}}$, we apply Lemma K.1 with $f$ being $(\gamma,\beta)\mapsto L(h_{\gamma},\beta)$, $x^{*}$ being $(h^{*},\beta^{*})$, $x^{\prime}$ being $(\hat{h},\hat{\beta})$, $\varepsilon=\mathcal{E}_{\textnormal{gen}}$, and the directions specified above, with their respective smoothness parameters derived above. Using the formulas for $\nabla L$ from Lemma N.5 and the bound $\|r_{\mathcal{H}_{B}}\|\leqslant B$, we obtain that on the event $E$,
$|2\lambda\mathbb{E}_{1}[\hat{h}(\hat{\beta}\hat{h}-r)]|\leqslant\mathcal{E}_{1},\qquad|\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]+\lambda\mathbb{E}_{1}[2\beta r_{\mathcal{H}_{B}}(\hat{\beta}\hat{h}-r)]|\leqslant\mathcal{E}_{2},$
@@
$\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]$ $=(\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]+\lambda\mathbb{E}_{1}[2\beta r_{\mathcal{H}_{B}}(\beta h-r)])-\lambda\mathbb{E}_{1}[2\beta(\beta h)(\beta h-r)]-\lambda\mathbb{E}_{1}[2\beta(r_{\mathcal{H}_{B}}-\beta h)(\beta h-r)].$
Evaluating at $(\hat{h},\hat{\beta})$, the first term is at most $\mathcal{E}_{2}$ in magnitude, the second term is at most $\hat{\beta}^{2}\mathcal{E}_{1}$ in magnitude, and the third term equals $2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]$. We deduce
$\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]\geqslant 2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]-\hat{\beta}^{2}\mathcal{E}_{1}-\mathcal{E}_{2}.$
@@
$|\mathbb{E}_{1}[r\widehat{\textnormal{Cover}}]-\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]|\leqslant(1-\alpha)\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|].$
@@
$\mathbb{E}_{1}[r\widehat{\textnormal{Cover}}]$ $\geqslant 2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]-\hat{\beta}^{2}\mathcal{E}_{1}-\mathcal{E}_{2}-(1-\alpha)\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|].$
@@
$A_{10}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/4}\lambda+A_{11}\left(\frac{1}{n_{1}^{1/4}}+\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/4}\right)\lambda^{1/2}+A_{12}\frac{1}{n_{1}^{1/4}}\lambda^{1/4}+A_{13}\frac{1}{n_{1}^{1/4}}+A_{14}\frac{1}{n_{1}^{1/4}}\lambda^{-1/4}$
and where $A_{10},\ldots,A_{14}$ are the positive constants given in Appendix F. It follows that on the event $E$,
$\mathbb{E}_{1}[r\widehat{\textnormal{Cover}}]\geqslant(1-\alpha)+2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]-\mathcal{E}_{\textnormal{cov}}-(1-\alpha)\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|].$
@@
Appendix L Unconstrained existence and boundedness
@@
In Lemma L.1, we show that under 10, we may eliminate $\beta$ from Equation LR-QR, so that Equation LR-QR is equivalent to solving the following unconstrained optimization problem over $h$:
@@
Lemma L.3 (Existence of unconstrained minimizers).
@@
Fix $\lambda\geqslant 0$. By 10 and Lemma L.1, it suffices to show that there exists a global minimizer of the objective in Equation 21. Let $G(h)$ denote the objective of Equation 21. Define the function $\tilde{h}=\theta^{*}r_{\mathcal{H}}\in\mathcal{H}\setminus\{0\}$, where $\theta^{*}$ is chosen to satisfy Lemma L.2. With $c_{\textnormal{indep}}$ from 6, define $\tilde{B}(\lambda):=2c_{\textnormal{indep}}^{-1}(1+\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\tilde{h},S)])>0$ and
@@
To see this, first suppose $\|h\|\geqslant\tilde{B}(\lambda)$. Then writing $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$ and applying Lemma N.1, the triangle inequality, and $S\in[0,1]$,
@@
Thus by Lemma N.1 and 8, $G(h)\geqslant\alpha(\tilde{B}(\lambda)c_{\textnormal{indep}}-1)-\lambda\mathbb{E}_{1}[r_{\mathcal{H}}^{2}]$. To prove the inequality $G(h)>G(\tilde{h})$, it suffices to show that
@@
Next, suppose $\|h\|\leqslant\tilde{b}(\lambda)$. By Lemma N.1, the triangle inequality, and $S\in[0,1]$,
@@
Lemma L.4 (Bounds on unconstrained minimizers).
Under the conditions used in Lemma L.3, for all $\lambda>0$, for any minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective in Equation LR-QR, we have that $\|h_{\lambda}^{*}\|\in(B_{\textnormal{lower}},B_{\textnormal{upper}})$ and $\beta_{\lambda}^{*}\in(\beta_{\textnormal{lower}},\beta_{\textnormal{upper}})$, where
@@
$B_{\textnormal{upper}}$ $=2c_{\textnormal{indep}}^{-1}(\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\theta^{*}r_{\mathcal{H}},S)]+1),\qquad\beta_{\textnormal{lower}}=\frac{c_{\textnormal{align}}}{B_{\textnormal{upper}}\lambda_{\textnormal{max}}(\Sigma)^{1/2}}>0,\qquad\beta_{\textnormal{upper}}=\frac{\mathbb{E}_{1}[r^{2}]^{1/2}}{B_{\textnormal{lower}}\lambda_{\textnormal{min}}(\Sigma)^{1/2}},$
and where $\theta^{*}>0$ is as in Lemma L.2 and $r_{\mathcal{H}}$ denotes the projection of $r$ onto $\mathcal{H}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$.
@@
for $\xi\geqslant 0$. We claim that for $\xi>0$, any minimizer of the objective in Equation 25 is of the form $h_{1/\xi}^{*}$. To see this, note that for $\xi>0$, the objective of Equation 21 with regularization $\lambda=1/\xi$ can be obtained by scaling the objective of Equation 25 by the positive factor $1/\xi$. Next, by 10, we may apply Lemma L.1 to deduce that $h\in\mathcal{H}\setminus\{0\}$ is a minimizer of the objective in Equation 21 with regularization $\lambda=1/\xi$ iff $h=h_{1/\xi}^{*}$.
In particular, by Lemma L.3, for all $\xi>0$, there exists a global minimizer of Equation 25 with regularization $\xi$. In the case that $\xi=0$, it is clear that any minimizer $h^{*}_{\infty}$ of the objective in Equation 25 with regularization $\xi=0$ has the form $h^{*}_{\infty}=\theta r_{\mathcal{H}}$ for some scalar $\theta>0$.
Since there exists a minimizer of the objective in Equation 25 for all regularizations $\xi$ in the interval $[0,\infty)$, we may apply Lemma M.1 to deduce that for all $\xi>0$ we have $\mathbb{E}_{1}[\ell_{\alpha}(h_{1/\xi}^{*},S)]\leqslant\mathbb{E}_{1}[\ell_{\alpha}(h_{\infty}^{*},S)]$.
@@
By Lemma L.2, there exists $\theta^{*}>0$ such that $\mathbb{E}_{1}[S]-\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\theta^{*}r_{\mathcal{H}},S)]>0$. Setting $h_{\infty}^{*}=\theta^{*}r_{\mathcal{H}}$ and plugging in the expression for $B_{\textnormal{lower}}$ given in (24), our lower bound becomes $\mathbb{E}_{1}[|h_{1/\xi}^{*}|]>\lambda_{\textnormal{max}}(\Sigma)^{1/2}B_{\textnormal{lower}}$. We now convert this $L^{1}$ norm bound to an $L^{2}$ norm bound as follows. Write $h_{1/\xi}^{*}=\langle\gamma_{1/\xi}^{*},\Phi\rangle$ for $\gamma_{1/\xi}^{*}\in\mathbb{R}^{d}$. By the Cauchy-Schwarz inequality, we obtain the upper bound
@@
Having established $0<B_{\textnormal{lower}}<\inf_{\lambda>0}\|h_{\lambda}^{*}\|\leqslant\sup_{\lambda>0}\|h_{\lambda}^{*}\|<B_{\textnormal{upper}}<\infty$, we turn to upper and lower bounds on $\beta^{*}_{\lambda}$. As shown in the proof of Lemma L.1, if $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ is a minimizer of the objective in Equation LR-QR with regularization $\lambda$, then $\beta_{\lambda}^{*}=\frac{\mathbb{E}_{1}[rh_{\lambda}^{*}]}{\mathbb{E}_{1}[|h_{\lambda}^{*}|^{2}]}$. By 7, $\frac{\mathbb{E}_{1}[rh_{0}^{*}]}{\mathbb{E}_{1}[|h_{0}^{*}|^{2}]^{1/2}}\geqslant c_{\textnormal{align}}>0$ for some minimizer $(h_{0}^{*},\beta_{0}^{*})$ of the objective in Equation LR-QR with regularization 00. By 10 and Lemma L.1, $h$ is a minimizer of the objective in Equation 21 with regularization $\lambda\geqslant 0$ iff $h=h_{\lambda}^{*}$ for some minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective in Equation LR-QR. Thus by Lemma L.3, for all $\lambda\geqslant 0$, there exists a global minimizer of Equation 21, and we may apply Lemma M.1 to Equation 21 to deduce that for any $\lambda\geqslant 0$ we have $\frac{\mathbb{E}_{1}[rh_{\lambda}^{*}]}{\mathbb{E}_{1}[|h_{\lambda}^{*}|^{2}]^{1/2}}\geqslant c_{\textnormal{align}}>0$. Consequently, by our bounds on $h_{\lambda}^{*}$, 8, and the Cauchy-Schwarz inequality, if we write $h_{\lambda}^{*}=\langle\gamma_{\lambda}^{*},\Phi\rangle$ for $\gamma_{\lambda}^{*}\in\mathbb{R}^{d}$, then we have
@@
Appendix M Monotonicity
Lemma M.1.
@@
Appendix N Helper lemmas
Lemma N.1.
@@
$0\leqslant\ell_{\alpha}(c_{1},s)-\ell_{\alpha}(c_{2},s)=(1-\alpha)(s-c_{1})-\alpha(c_{2}-s)\leqslant(1-\alpha)(s-c_{1})+(1-\alpha)(c_{2}-s)=(1-\alpha)(c_{2}-c_{1}),$
@@
$=\mathbb{E}_{1}[(\mathbb{P}_{S|X}[h(X)>S]-(1-\alpha))\langle\tilde{\gamma},\Phi(X)\rangle]=\mathbb{E}_{1}[(\mathbb{P}_{S|X}[h(X)>S]-(1-\alpha))g(X)].$

### v2

@@ Appendix I Lipschitz process
By Lemma M.3 and Lemma M.4, a global minimizer of the objective in Equation LR-QR exists, and since $\beta_{\textnormal{min}}\leqslant\beta_{\textnormal{lower}}$, $\beta_{\textnormal{max}}\geqslant\beta_{\textnormal{upper}}$, and $B\geqslant B_{\textnormal{upper}}$, any such minimizer lies in the interior of $\mathcal{H}_{B}\times\mathcal{I}$. Thus we may apply Lemma L.1 to the objective function $L$. We utilize two directional derivatives in the space $\mathcal{H}\times\mathbb{R}$. The first is in the direction $0_{\mathcal{H}}\times 1$, the unit vector in the $\beta$ coordinate. Since $(\hat{h},\hat{\beta})\in\mathcal{H}_{B}\times\mathcal{I}$, the magnitude of the second derivative of $L$ along this direction is bounded by $\nu_{1}$.
The second is in the direction of the vector $r_{B}\times 0$, where $r_{B}$ the projection of $r$ onto the closed convex set $\mathcal{H}_{B}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$. Since $(\hat{h},\hat{\beta})\in\mathcal{H}_{B}\times\mathcal{I}$, the magnitude of the second derivative of $L$ along this direction is bounded by $\nu_{2}$.
Given $\hat{h}$, let $\widehat{\textnormal{Cover}}(X):=\mathbb{P}\left[S\leqslant\hat{h}(X)|X\right]-(1-\alpha)$. Now, on the event $E$ that $L(\hat{h},\hat{\beta})-L(h^{*},\beta^{*})\leqslant\mathcal{E}_{\textnormal{gen}}$, we apply Lemma L.1 with $f$ being $(\gamma,\beta)\mapsto L(h_{\gamma},\beta)$, $x^{*}$ being $(h^{*},\beta^{*})$, $x^{\prime}$ being $(\hat{h},\hat{\beta})$, $\varepsilon=\mathcal{E}_{\textnormal{gen}}$, and the directions specified above, with their respective smoothness parameters derived above. Using the formulas for $\nabla L$ from Lemma O.5 and the bound $\|r_{B}\|\leqslant B$, we obtain that on the event $E$,
$|2\lambda\mathbb{E}_{1}[\hat{h}(\hat{\beta}\hat{h}-r)]|\leqslant\mathcal{E}_{1},\qquad|\mathbb{E}_{1}[r_{B}\widehat{\textnormal{Cover}}]+\lambda\mathbb{E}_{1}[2\beta r_{B}(\hat{\beta}\hat{h}-r)]|\leqslant\mathcal{E}_{2},$
@@
$\mathbb{E}_{1}[r_{B}\widehat{\textnormal{Cover}}]$ $=(\mathbb{E}_{1}[r_{B}\widehat{\textnormal{Cover}}]+\lambda\mathbb{E}_{1}[2\beta r_{B}(\beta h-r)])$
$-\lambda\mathbb{E}_{1}[2\beta(\beta h)(\beta h-r)]-\lambda\mathbb{E}_{1}[2\beta(r_{B}-\beta h)(\beta h-r)].$
Evaluating at $(\hat{h},\hat{\beta})$, the first term is at most $\mathcal{E}_{2}$ in magnitude, the second term is at most $\hat{\beta}^{2}\mathcal{E}_{1}$ in magnitude, and the third term equals $2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{B}-\hat{\beta}\hat{h})^{2}]$. We deduce
$\mathbb{E}_{1}[r_{B}\widehat{\textnormal{Cover}}]\geqslant 2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{B}-\hat{\beta}\hat{h})^{2}]-\hat{\beta}^{2}\mathcal{E}_{1}-\mathcal{E}_{2}.$
@@
$|\mathbb{E}_{1}[r\widehat{\textnormal{Cover}}]-\mathbb{E}_{1}[r_{B}\widehat{\textnormal{Cover}}]|\leqslant(1-\alpha)\mathbb{E}_{1}[|r-r_{B}|].$
@@
$\mathbb{E}_{1}[r\widehat{\textnormal{Cover}}]$ $\geqslant 2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{B}-\hat{\beta}\hat{h})^{2}]-\hat{\beta}^{2}\mathcal{E}_{1}-\mathcal{E}_{2}-(1-\alpha)\mathbb{E}_{1}[|r-r_{B}|].$
@@
$A_{10}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/4}\lambda+A_{11}\left(\frac{1}{n_{1}^{1/4}}+\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/4}\right)\lambda^{1/2}+A_{12}\frac{\lambda^{1/4}}{n_{1}^{1/4}}+\frac{A_{13}}{n_{1}^{1/4}}+A_{14}\frac{\lambda^{-1/4}}{n_{1}^{1/4}}$
and where $A_{10},\ldots,A_{14}$ are the positive constants given in Appendix G. It follows that on the event $E$,
$\mathbb{E}_{1}[r\widehat{\textnormal{Cover}}]\geqslant(1-\alpha)+2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{B}-\hat{\beta}\hat{h})^{2}]-\mathcal{E}_{\textnormal{cov}}-(1-\alpha)\mathbb{E}_{1}[|r-r_{B}|].$
@@
Appendix M Unconstrained existence and boundedness
@@
In Lemma M.1, we show that under 10, we may eliminate $\beta$ from Equation LR-QR, so that Equation LR-QR is equivalent to solving the following unconstrained optimization problem over $h$:
@@
Lemma M.3 (Existence of unconstrained minimizers).
@@
Fix $\lambda\geqslant 0$. By 10 and Lemma M.1, it suffices to show that there exists a global minimizer of the objective in Equation 19. Let $G(h)$ denote the objective of Equation 19. Define the function $\tilde{h}=\theta^{*}r_{\mathcal{H}}\in\mathcal{H}\setminus\{0\}$, where $\theta^{*}$ is chosen to satisfy Lemma M.2. With $c_{\textnormal{indep}}$ from 6, define $\tilde{B}(\lambda):=2c_{\textnormal{indep}}^{-1}(1+\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\tilde{h},S)])>0$ and
@@
To see this, first suppose $\|h\|\geqslant\tilde{B}(\lambda)$. Then writing $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$ and applying Lemma O.1, the triangle inequality, and $S\in[0,1]$,
@@
Thus by Lemma O.1 and 8, $G(h)\geqslant\alpha(\tilde{B}(\lambda)c_{\textnormal{indep}}-1)-\lambda\mathbb{E}_{1}[r_{\mathcal{H}}^{2}]$. To prove the inequality $G(h)>G(\tilde{h})$, it suffices to show that
@@
Next, suppose $\|h\|\leqslant\tilde{b}(\lambda)$. By Lemma O.1, the triangle inequality, and $S\in[0,1]$,
@@
Lemma M.4 (Bounds on unconstrained minimizers).
Under the conditions used in Lemma M.3, for all $\lambda>0$, for any minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective in Equation LR-QR, we have that $\|h_{\lambda}^{*}\|\in(B_{\textnormal{lower}},B_{\textnormal{upper}})$ and $\beta_{\lambda}^{*}\in(\beta_{\textnormal{lower}},\beta_{\textnormal{upper}})$, where
@@
$B_{\textnormal{upper}}$ $=2c_{\textnormal{indep}}^{-1}(\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\theta^{*}r_{\mathcal{H}},S)]+1),\qquad\beta_{\textnormal{lower}}=\frac{c_{\textnormal{align}}}{B_{\textnormal{upper}}\lambda_{\textnormal{max}}(\Sigma)^{1/2}}>0,$
$\beta_{\textnormal{upper}}$ $=\frac{\mathbb{E}_{1}[r^{2}]^{1/2}}{B_{\textnormal{lower}}\lambda_{\textnormal{min}}(\Sigma)^{1/2}},$
and where $\theta^{*}>0$ is as in Lemma M.2 and $r_{\mathcal{H}}$ denotes the projection of $r$ onto $\mathcal{H}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$.
@@
for $\xi\geqslant 0$. We claim that for $\xi>0$, any minimizer of the objective in Equation 23 is of the form $h_{1/\xi}^{*}$. To see this, note that for $\xi>0$, the objective of Equation 19 with regularization $\lambda=1/\xi$ can be obtained by scaling the objective of Equation 23 by the positive factor $1/\xi$. Next, by 10, we may apply Lemma M.1 to deduce that $h\in\mathcal{H}\setminus\{0\}$ is a minimizer of the objective in Equation 19 with regularization $\lambda=1/\xi$ iff $h=h_{1/\xi}^{*}$.
In particular, by Lemma M.3, for all $\xi>0$, there exists a global minimizer of Equation 23 with regularization $\xi$. In the case that $\xi=0$, it is clear that any minimizer $h^{*}_{\infty}$ of the objective in Equation 23 with regularization $\xi=0$ has the form $h^{*}_{\infty}=\theta r_{\mathcal{H}}$ for some scalar $\theta>0$.
Since there exists a minimizer of the objective in Equation 23 for all regularizations $\xi$ in the interval $[0,\infty)$, we may apply Lemma N.1 to deduce that for all $\xi>0$ we have $\mathbb{E}_{1}[\ell_{\alpha}(h_{1/\xi}^{*},S)]\leqslant\mathbb{E}_{1}[\ell_{\alpha}(h_{\infty}^{*},S)]$.
@@
By Lemma M.2, there exists $\theta^{*}>0$ such that $\mathbb{E}_{1}[S]-\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\theta^{*}r_{\mathcal{H}},S)]>0$. Setting $h_{\infty}^{*}=\theta^{*}r_{\mathcal{H}}$ and plugging in the expression for $B_{\textnormal{lower}}$ given in (22), our lower bound becomes $\mathbb{E}_{1}[|h_{1/\xi}^{*}|]>\lambda_{\textnormal{max}}(\Sigma)^{1/2}B_{\textnormal{lower}}$. We now convert this $L^{1}$ norm bound to an $L^{2}$ norm bound as follows. Write $h_{1/\xi}^{*}=\langle\gamma_{1/\xi}^{*},\Phi\rangle$ for $\gamma_{1/\xi}^{*}\in\mathbb{R}^{d}$. By the Cauchy-Schwarz inequality, we obtain the upper bound
@@
Having established $0<B_{\textnormal{lower}}<\inf_{\lambda>0}\|h_{\lambda}^{*}\|\leqslant\sup_{\lambda>0}\|h_{\lambda}^{*}\|<B_{\textnormal{upper}}<\infty$, we turn to upper and lower bounds on $\beta^{*}_{\lambda}$. As shown in the proof of Lemma M.1, if $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ is a minimizer of the objective in Equation LR-QR with regularization $\lambda$, then $\beta_{\lambda}^{*}=\frac{\mathbb{E}_{1}[rh_{\lambda}^{*}]}{\mathbb{E}_{1}[|h_{\lambda}^{*}|^{2}]}$. By 7, $\frac{\mathbb{E}_{1}[rh_{0}^{*}]}{\mathbb{E}_{1}[|h_{0}^{*}|^{2}]^{1/2}}\geqslant c_{\textnormal{align}}>0$ for some minimizer $(h_{0}^{*},\beta_{0}^{*})$ of the objective in Equation LR-QR with regularization 00. By 10 and Lemma M.1, $h$ is a minimizer of the objective in Equation 19 with regularization $\lambda\geqslant 0$ iff $h=h_{\lambda}^{*}$ for some minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective in Equation LR-QR. Thus by Lemma M.3, for all $\lambda\geqslant 0$, there exists a global minimizer of Equation 19, and we may apply Lemma N.1 to Equation 19 to deduce that for any $\lambda\geqslant 0$ we have $\frac{\mathbb{E}_{1}[rh_{\lambda}^{*}]}{\mathbb{E}_{1}[|h_{\lambda}^{*}|^{2}]^{1/2}}\geqslant c_{\textnormal{align}}>0$. Consequently, by our bounds on $h_{\lambda}^{*}$, 8, and the Cauchy-Schwarz inequality, if we write $h_{\lambda}^{*}=\langle\gamma_{\lambda}^{*},\Phi\rangle$ for $\gamma_{\lambda}^{*}\in\mathbb{R}^{d}$, then we have
@@
Appendix N Monotonicity
Lemma N.1.
@@
Appendix O Helper lemmas
Lemma O.1.
@@
$0\leqslant\ell_{\alpha}(c_{1},s)-\ell_{\alpha}(c_{2},s)=(1-\alpha)(s-c_{1})-\alpha(c_{2}-s)$
$\leqslant(1-\alpha)(s-c_{1})+(1-\alpha)(c_{2}-s)=(1-\alpha)(c_{2}-c_{1}),$
@@
$=\mathbb{E}_{1}[(\mathbb{P}_{S|X}[h(X)>S]-(1-\alpha))\langle\tilde{\gamma},\Phi(X)\rangle]$
$=\mathbb{E}_{1}[(\mathbb{P}_{S|X}[h(X)>S]-(1-\alpha))g(X)].$